OpenStudy (anonymous):

Please solve! I need to check my work 2x+5 x+1 ---------- + ----- x^2-3x-10 x+2

OpenStudy (johnweldon1993):

\[\large \frac{2x + 5}{(x - 5)(x + 2)} + \frac{x + 1}{x + 2}\] what will be the common denominator?

OpenStudy (johnweldon1993):

It looks like if we multiply the second fraction by (x - 5) we will have the same denominator as the first fraction... \[\large \frac{2x + 5}{(x - 5)(x + 2)} + \frac{(x + 1)(x - 5)}{(x -5)(x + 2)}\]

OpenStudy (johnweldon1993):

Now we can combine these fractions over the common denominator \[\large \frac{2x + 5 + (x + 1)(x - 5)}{(x - 5)(x + 2)}\]

OpenStudy (johnweldon1993):

Lets distribute out that quadratic there \[\large \frac{2x + 5 + x^2 - 4x - 5}{(x - 5)(x + 2)}\] Simplify that a bit \[\large \frac{x^2 - 2x}{(x - 5)(x + 2)}\] And lets just factor out an 'x' from the top there \[\large \frac{x(x - 2)}{(x - 5)(x + 2)}\] This will be our simplified answer

OpenStudy (anonymous):

okay thats what I got. thank you!

OpenStudy (johnweldon1993):

Anytime :)